# Antisymmetric tensor influence on charged black hole lensing phenomena and time delay

### v1

@@ Abstract
In this work, we investigate the gravitational lensing of a charged black hole (spherically symmetric) within the context of Lorentz violation triggered by an antisymmetric tensor field. Our calculations consider two different scenarios: the weak and strong deflection limits. For the weak deflection limit, we employ the Gauss–Bonnet theorem to obtain our results. In the strong deflection limit, we utilize the Tsukamoto methodology, which provides measurable outcomes such as relativistic image positions and magnifications. Applying the latter methodology, we analyze the gravitational lensing by Sagittarius $A^{*}$ and derive the related observables, which are expressed as functions of the Lorentz violation parameter. Finally, the time delay is addrressed as well.
@@ I Introduction
Beyond the conventional vector field theories, an alternative approach to investigating Lorentz symmetry breaking (LSB) involves the study of a rank–two antisymmetric tensor field known as the Kalb–Ramond field 042 ; assunccao2019dynamical ; maluf2019antisymmetric . This field is inherently present in the spectrum of bosonic string theory 043 . When it is non–minimally coupled to gravity and acquires a non–zero vacuum expectation value, spontaneous breaking of Lorentz symmetry is induced. An exact solution for a static, spherically symmetric configuration within this framework has been presented by Ref. 044 . Subsequently, an extensive investigation into the dynamics of both massive and massless particles near such static spherical Kalb–Ramond black holes has been conducted by Ref. 045 . Furthermore, the gravitational deflection of light and the shadows cast by rotating black holes within this theoretical framework have been explored by Ref. 046 . Additionally, recent literature has addressed the Generalized Uncertainty Principle (GUP) corrections, gravitational parity violations involving antisymmetric tensors, and the cosmological implications of Kalb–Ramond–like particles baruah2023quasinormal ; manton2024kalb ; capanelli2023cosmological .
@@
Recent observations made by the Event Horizon Telescope of a supermassive black hole situated at the heart of the M87 galaxy have generated profound scientific interest 023 ; 024 ; 025 ; 026 ; 027 ; 028 ; 029 . The foundational work by Virbhadra and Ellis introduced a concise lens equation tailored for supermassive black holes within an asymptotically flat background 030 ; 031 , revealing multiple symmetrically distributed images around the optical axis due to strong gravitational effects. Subsequent advancements spearheaded by Fritelli et al. 032 , Bozza et al. 033 , and Tsukamoto 035 have refined the analytical frameworks for investigating strong field gravitational lensing. These studies have explored light deflection across various scenarios grespan2023strong ; cunha2018shadows ; oguri2019strong ; metcalf2019strong ; bisnovatyi2017gravitational ; ezquiaga2021phase , including exotic structures such as wormholes ovgun2019exact ; 38.1 ; 38.2 ; 38.3 ; 38.4 ; 38.5 , rotating solutions jusufi2018gravitational ; 37.1 ; 37.2 ; 37.3 ; 37.4 ; 37.5 ; 37.6 , alternative theories of gravity chakraborty2017strong ; t25:1 ; 40 ; nascimento2024gravitational , and Reissner–Nordström spacetime 036 ; 036.1 ; 036.2 , and others zhang2024strong ; tsukamoto2023gravitational .
@@ III Gravitational lensing via weak deflection limit
In this section, we conduct an examination of the Gauss–Bonnet theorem and proceed to calculate the weak deflection angle of the black hole. Initially, we derive the null geodesics under the condition $\mathrm{d}s^{2}=0$, which upon rearrangement gives:
$\mathrm{d}t^{2}=\gamma_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}=\frac{1}{f(r)^{2}}\mathrm{d}r^{2}+\frac{r^{2}}{f(r)}\mathrm{d}\Omega^{2},\leavevmode\nobreak$ (2)
Here, indices $i$ and $j$ range from $1$ to $3$, and $\gamma_{ij}$ denotes the optical metric. Utilizing the Gauss–Bonnet theorem requires us to calculate the Gaussian curvature, which is carried out as follows:
$\begin{split}\mathcal{K}&=\frac{R}{2}=\frac{f(r)}{2}\frac{\mathrm{d}^{2}}{\mathrm{d}r^{2}}f(r)-\frac{\left(\frac{\mathrm{d}}{\mathrm{d}r}f(r)\right)^{2}}{4}\\ &=-\frac{6MQ^{2}}{(1-l)^{2}r^{5}}-\frac{2M}{(1-l)r^{3}}+\frac{2Q^{4}}{(1-l)^{4}r^{6}}+\frac{3Q^{2}}{(1-l)^{3}r^{4}}+\frac{3M^{2}}{r^{4}}.\end{split}$ (3)
Here, $\gamma$ denotes the determinant of $\gamma_{ij}$, and $R$ stands for the Ricci scalar. The area of the surface on the equatorial plane is given by Gibbons:2008rj :
$\mathrm{d}S=\sqrt{\gamma}\mathrm{d}r\mathrm{d}\phi=\frac{r}{f(r)^{3/2}}\mathrm{d}r\mathrm{d}\phi=\left(\frac{r}{\left(\frac{Q^{2}}{(1-l)^{2}r^{2}}+\frac{1}{1-l}-\frac{2M}{r}\right)^{3/2}}\right)\mathrm{d}r\mathrm{d}\phi.$ (4)
@@
$\begin{split}\alpha(b,l,Q)&=-\int\int\mathcal{K}\mathrm{d}S=-\int^{\pi}_{0}\int^{\infty}_{\frac{b}{\sin\phi}}\mathcal{K}\mathrm{d}S\\ \simeq&\frac{4M}{b}+\frac{3\pi M^{2}}{4b^{2}}-\frac{2lM}{b}-\frac{9\pi lM^{2}}{8b^{2}}-\frac{9\pi lQ^{2}}{8b^{2}}-\frac{3\pi Q^{2}}{4b^{2}}\\ &-\frac{8MQ^{2}}{3b^{3}}-\frac{4lMQ^{2}}{3b^{3}}+\frac{45\pi lM^{2}Q^{2}}{64b^{4}}-\frac{45\pi M^{2}Q^{2}}{32b^{4}}.\end{split}$ (5)
@@ IV Gravitational lensing via the strong deflection limit
This section aims to outline the general methodology described in tsukamoto2017deflection , which we will use to calculate the deflection angle of a light ray in the strong deflection limit. In our analysis, similar to many prior studies (e.g., Nascimento:2020ime ), we focus on asymptotically flat, static, and spherically symmetric spacetimes, characterized by the following line element $\mathrm{d}s^{2}=-A(r)\mathrm{d}t^{2}+B(r)\mathrm{d}r^{2}+C(r)(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2})$. To use the general methodology addressed by Tsukamoto tsukamoto2017deflection , we need the metric to match the condition of asymptotic flatness. Due to the symmetries of the spacetime, there are two Killing vectors: $\partial_{t}$ and $\partial_{\phi}$, as one should expect.
Let us now outline the procedure for calculating the deflection angle in the strong field regime. Initially, we introduce a new variable denoted as $D(r)$: $\equiv\frac{C^{\prime}(r)}{C(r)}-\frac{A^{\prime}(r)}{A(r)}$. With these definitions, we introduce the impact parameter $b$ as follows: $b\equiv\frac{L}{E}=\frac{C(r)\dot{\phi}}{A(r)\dot{t}}$. Due to the axial symmetry, the motion can be confined to the equatorial plane, where $\theta=\pi/2$. Consequently, the radial equation simplifies to $\dot{r}^{2}=V(r)$. The potential function is given by $V(r)=\frac{L^{2}R(r)}{B(r)C(r)}$, where $R(r)\equiv\frac{C(r)}{A(r)}\frac{b^{2}}{b^{2}-1}$. Consequently, $V(r)$ diminishes at $r=r_{0}$. From the trajectory equation, we conclude that $A_{0}\dot{t}^{2}_{0}=C_{0}\dot{\phi}^{2}_{0}$.
In this context and throughout subsequent discussions, the subscript ``$0"$ signifies quantities evaluated at $r=r_{0}$. For simplification without compromising generality, we assume the impact parameter $b$ to be positive, particularly when dealing with a single light ray. Due to the preservation of the impact parameter along the trajectory, which can be expressed as $b(r_{0})=\frac{L}{E}=\frac{C_{0}\dot{\phi}_{0}}{A_{0}\dot{t}_{0}}=\sqrt{\frac{C_{0}}{A_{0}}}$. Accordingly, $R(r)$ can be expressed in the following manner $R(r)=\frac{A_{0}C}{AC_{0}}-1$.
Therefore, the trajectory equation takes the following form $\left(\frac{\mathrm{d}r}{\mathrm{d}\phi}\right)^{2}=\frac{R(r)C(r)}{B(r)}$, and, thereby, we can address the deflection angle of light, denoted as $\alpha(r_{0})$, which is given $\alpha(r_{0})=I(r_{0})-\pi$, where, $I(r_{0})$ is properly defined as being $I(r_{0})\equiv 2\int^{\infty}_{r_{0}}\frac{\mathrm{d}r}{\sqrt{\frac{R(r)C(r)}{B(r)}}}$. The critical step in our calculation now involves evaluating this integral, which can be a challenging task depending on the metric under consideration tsukamoto2017deflection . For this regard, to find $I(r_{0})$, let us introduce tsukamoto2017deflection $z\equiv 1-\frac{r_{0}}{r}$, so that $I(r_{0})=\int^{1}_{0}f(z,r_{0})\mathrm{d}z$ in which $f(z,z_{0})\equiv\frac{2r_{0}}{\sqrt{G(z,r_{0})}},\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,G(z,r_{0})\equiv R\frac{C}{B}(1-z)^{4}$. Thereby, $G(z,r_{0})$ can be succinctly formulated tsukamoto2017deflection $G_{m}(z)=c_{2}(r_{m})z^{2}+\mathcal{O}(z^{3})$. This indicates that the principal divergence of $f(z,r_{0})$ arises at the $z^{-1}$ order, resulting in a logarithmic divergence in the integral $I(r_{0})$ as we approach the strong deflection limit $r_{0}\to r_{m}$. To handle this divergence, we decompose the integral $I(r_{0})$ into two components: a divergent segment $I_{D}(r_{0})$ and a well–behaved segment $I_{R}(r_{0})$. The divergent part $I_{D}(r_{0})$ is defined as follows $I_{D}(r_{0})\equiv\int^{1}_{0}f_{D}(z,r_{0})\mathrm{d}z,\text{with}\,f_{D}(z,r_{0})\equiv\frac{2r_{0}}{\sqrt{c_{1}(r_{0})z+c_{2}(r_{0})z^{2}}}$.
To proceed forward, the divergent part can be computed $I_{D}(b)=-\frac{r_{m}}{\sqrt{c_{2}(r_{m})}}\ln\left[\frac{b}{b_{c}}-1\right]+\frac{r_{m}}{\sqrt{c_{2}(r_{m})}}\ln\left[r^{2}D^{\prime}_{m}\right]+\mathcal{O}[(b-b_{c})\ln(b-b_{c})]$. In contrast, it convenient to define the regular part as $I_{R}(b)=\int^{0}_{1}f_{R}(z,b_{c})\mathrm{d}z+\mathcal{O}[(b-b_{c})\ln(b-b_{c})]$, with $f_{R}\equiv f(z,r_{0})-f_{D}(z,r_{0})$ and the respective deflection angle is given by $a(b)=-\tilde{a}\ln\left[\frac{b}{b_{c}}-1\right]+\tilde{b}+\mathcal{O}[(b-b_{c})\ln(b-b_{c})]$, where $\tilde{a}=\sqrt{\frac{2B_{m}A_{m}}{C^{\prime\prime}_{m}A_{m}-C_{m}A^{\prime\prime}_{m}}},\text{and}\tilde{b}=\tilde{a}\ln\left[r^{2}_{m}\left(\frac{C^{\prime\prime}}{C_{m}}-\frac{A^{\prime\prime}_{m}}{C_{m}}\right)\right]+I_{R}(r_{m})-\pi.$

### v2

@@ Abstract
In this work, we investigate the gravitational lensing of a charged black hole (spherically symmetric) within the context of Lorentz violation triggered by an antisymmetric tensor field. Our calculations consider two different scenarios: the weak and strong deflection limits. For the weak deflection limit, we employ the Gauss–Bonnet theorem to obtain our results. In the strong deflection limit, we utilize the Tsukamoto methodology, which provides measurable outcomes such as relativistic image positions and magnifications. Applying the latter methodology, we analyze the gravitational lensing by Sagittarius $A^{*}$ and derive the related observables, which are expressed as functions of the Lorentz violation parameter. Finally, the time delay is addressed as well.
@@ I Introduction
Beyond the conventional vector field theories, an alternative approach to investigating Lorentz symmetry breaking (LSB) involves the study of a rank–two antisymmetric tensor field known as the Kalb–Ramond field 042 ; assunccao2019dynamical ; maluf2019antisymmetric ; AraujoFilho:2024ctw . This field is inherently present in the spectrum of bosonic string theory 043 . When it is non–minimally coupled to gravity and acquires a non–zero vacuum expectation value, spontaneous breaking of Lorentz symmetry is induced. An exact solution for a static, spherically symmetric configuration within this framework has been presented by Ref. 044 . Subsequently, an extensive investigation into the dynamics of both massive and massless particles near such static spherical Kalb–Ramond black holes has been conducted by Ref. 045 . Furthermore, the gravitational deflection of light and the shadows cast by rotating black holes within this theoretical framework have been explored by Ref. 046 . Additionally, recent literature has addressed the Generalized Uncertainty Principle (GUP) corrections, gravitational parity violations involving antisymmetric tensors, and the cosmological implications of Kalb–Ramond–like particles baruah2023quasinormal ; manton2024kalb ; capanelli2023cosmological .
@@
Recent observations made by the Event Horizon Telescope of a supermassive black hole situated at the heart of the M87 galaxy have generated profound scientific interest 023 ; 024 ; 025 ; 026 ; 027 ; 028 ; 029 . The foundational work by Virbhadra and Ellis introduced a concise lens equation tailored for supermassive black holes within an asymptotically flat background 030 ; 031 , revealing multiple symmetrically distributed images around the optical axis due to strong gravitational effects. Subsequent advancements spearheaded by Fritelli et al. 032 , Bozza et al. 033 , and Tsukamoto 035 have refined the analytical frameworks for investigating strong field gravitational lensing. These studies have explored light deflection across various scenarios virbhadra2002gravitational ; virbhadra1998role ; virbhadra2000schwarzschild ; grespan2023strong ; cunha2018shadows ; oguri2019strong ; metcalf2019strong ; bisnovatyi2017gravitational ; ezquiaga2021phase , including exotic structures such as wormholes ovgun2019exact ; 38.1 ; 38.2 ; 38.3 ; 38.4 ; 38.5 , rotating solutions hsieh2021strong ; hsieh2021gravitational ; jusufi2018gravitational ; 37.1 ; 37.2 ; 37.3 ; 37.4 ; 37.5 ; 37.6 , alternative theories of gravity chakraborty2017strong ; t25:1 ; 40 ; nascimento2024gravitational , and Reissner–Nordström spacetime 036 ; 036.1 ; 036.2 , and others zhang2024strong ; tsukamoto2023gravitational . Also, gravitational distortion has also been addressed virbhadra2024conservation ; virbhadra2022distortions
@@ III Gravitational lensing via weak deflection limit
In this section, we conduct an examination of the Gauss–Bonnet theorem and proceed to calculate the weak deflection angle of the black hole Gibbons:2008rj . Adopting this framework, one can describe the trajectories of massless particles—those governed by the condition $\mathrm{d}s^{2}=0$—through the following expression, as detailed in Refs. araujo2024effects ; Heidari:2025iiv ; heidari2024absorption :
$\mathrm{d}t^{2}=\gamma_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}=-\frac{g_{rr}(r)}{g_{tt}(r)}\mathrm{d}r^{2}-\frac{\bar{g}_{\varphi\varphi}(r)}{g_{tt}(r)}\mathrm{d}\varphi^{2}.$ (2)
In this context, the indices $i$ and $j$ run over the spatial coordinates $1$ to $3$, and the quantity $\gamma_{ij}$ denotes the components of the optical metric. For convenience, the angular component evaluated on the equatorial plane is defined as $\bar{g}_{\varphi\varphi}(r)\equiv g_{\varphi\varphi}(r,\theta=\pi/2)$. Moreover, the Gaussian curvature associated with the optical geometry is expressed as qiao2024existence :
$\mathcal{K}(r,l,Q)=\frac{R}{2}=\frac{g_{tt}(r)}{\sqrt{g_{rr}(r)\,\bar{g}_{\varphi\varphi}(r)}}\frac{\partial}{\partial r}\left[\frac{g_{tt}(r)}{2\sqrt{g_{rr}(r)\,\bar{g}_{\varphi\varphi}(r)}}\frac{\partial}{\partial r}\left(\frac{\bar{g}_{\varphi\varphi}(r)}{g_{tt}(r)}\right)\right].$ (3)
Here, $R$ denotes the two–dimensional Ricci scalar associated with the optical geometry. In the regime where the parameter $l$ is small, the curvature can be approximated by the following expression heidari2024impact ; al-Badawi:2024pdx :
$\begin{split}&\mathcal{K}(r,l,Q)=+\frac{3Q^{2}}{(1-l)^{3}r^{4}}+\frac{3M^{2}}{r^{4}}-\frac{6MQ^{2}}{(1-l)^{2}r^{5}}-\frac{2M}{(1-l)r^{3}}+\frac{2Q^{4}}{(1-l)^{4}r^{6}}.\end{split}$ (4)
The area of the surface on the equatorial plane is given by qiao2024existence ; Heidari:2025iiv :
$\mathrm{d}S=\sqrt{\gamma}\,\mathrm{d}r\mathrm{d}\varphi=\sqrt{\frac{g_{rr}(r)}{g_{tt}(r)}\frac{g_{\varphi\varphi}(r)}{g_{tt}(r)}}\,\mathrm{d}r\mathrm{d}\varphi.$ (5)
@@
$\begin{split}\alpha(b,l,Q)&=-\int\int\mathcal{K}\mathrm{d}S=-\int^{\pi}_{0}\int^{\infty}_{r=\frac{b}{\sin\phi}}\mathcal{K}\mathrm{d}S\\ \simeq&\frac{4M}{b}+\frac{3\pi M^{2}}{4b^{2}}-\frac{2lM}{b}-\frac{9\pi lM^{2}}{8b^{2}}-\frac{9\pi lQ^{2}}{8b^{2}}-\frac{3\pi Q^{2}}{4b^{2}}\\ &-\frac{8MQ^{2}}{3b^{3}}-\frac{4lMQ^{2}}{3b^{3}}+\frac{45\pi lM^{2}Q^{2}}{64b^{4}}-\frac{45\pi M^{2}Q^{2}}{32b^{4}}.\end{split}$ (6)
@@
An additional remark worth considering is that it is not mandatory to adopt $b/\sin\phi$ as the lower integration limit in Eq. (6). Recent studies, for instance, have extended the analysis to include higher–order terms Jha:2024ltc ; AraujoFilho:2024mvz
$u=\frac{1}{r}=\frac{\sin\phi}{b}+\frac{M(1-\cos\phi)^{2}}{b^{2}}-\frac{M^{2}(60\phi\,\cos\phi+3\sin 3\phi-5\sin\phi)}{16b^{3}}.$ (7)
@@ IV Gravitational lensing via the strong deflection limit
The approach employed in this part of the analysis aims to derive the expression for the bending angle of light in the regime of strong gravitational deflection nascimento2024gravitational ; heidari2024absorption . Among the established frameworks available, we adopt the procedure initially developed by Tsukamoto tsukamoto2017deflection , which is particularly suited to probe photon trajectories in regions where the gravitational field is extremely intense. The spacetime geometry under consideration is defined by a general line element of the form:
$\mathrm{d}s^{2}=-\mathcal{A}(r)\mathrm{d}t^{2}+\mathcal{B}(r)\mathrm{d}r^{2}+\mathcal{C}(r)(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}).$ (8)
In order to implement the formalism established by Tsukamoto tsukamoto2017deflection , one must ensure that the underlying spacetime geometry adheres to the condition of asymptotic flatness in principle. This imposes specific constraints on the metric functions: as $r$ approaches infinity, the temporal and radial components must converge to unity, i.e., $\mathcal{A}(r)\to 1$ and $\mathcal{B}(r)\to 1$, while the angular part should asymptotically approach $\mathcal{C}(r)\to r^{2}$. The static and spherically symmetric nature of the geometry implies the existence of conserved quantities associated with the Killing vectors $\partial_{t}$ and $\partial_{\phi}$, which correspond to energy and angular momentum, respectively.
To advance toward the derivation of the deflection angle in the strong deflection regime, we begin by introducing an auxiliary function $\mathcal{D}(r)$, which simplifies the mathematical structure of the integral formulation used in the analysis
$\mathcal{D}(r)\equiv\frac{\mathcal{C}^{\prime}(r)}{\mathcal{C}(r)}-\frac{\mathcal{A}^{\prime}(r)}{\mathcal{A}(r)}.$ (9)
Derivatives with respect to the radial coordinate are denoted by the prime ($\prime$) symbol. It is presumed that the equation $\mathcal{D}(r)=0$ admits at least one positive root. Among these, the photon sphere is identified by selecting the greatest such root, which we label as $r_{ph}$. For the analysis to remain valid in the region extending from the photon sphere outward, the metric functions $\mathcal{A}(r)$, $\mathcal{B}(r)$, and $\mathcal{C}(r)$ are required to be well–defined, strictly positive, and finite for all $r\geq r_{ph}$.
The symmetries of the spacetime, expressed through the Killing vectors $\partial_{t}$ and $\partial_{\phi}$, lead to the conservation of two physical quantities along the null geodesics: the energy $E$ and the angular momentum $L$. These are given by $E=\mathcal{A}(r)\dot{t}$ and $L=\mathcal{C}(r)\dot{\phi}$, respectively. Assuming that neither $E$ nor $L$ vanishes, one can define the impact parameter $b$, which characterizes the trajectory of the photon, as follows:
$b\equiv\frac{L}{E}=\frac{\mathcal{C}(r)\dot{\phi}}{\mathcal{A}(r)\dot{t}}.$ (10)
Taking advantage of the axial symmetry of the spacetime, the analysis can be confined to the equatorial plane by setting $\theta=\pi/2$, without any loss of generality. Under this condition, the geodesic equation simplifies, allowing the radial component of the motion to be expressed as:
$\dot{r}^{2}=V(r).$ (11)
By defining the function $R(r)$ as $R(r)\equiv\mathcal{C}(r)/[\mathcal{A}(r)b^{2}]-1$, one can express the radial equation in terms of an effective potential, $V(r)$, given by $V(r)=L^{2}R(r)/[\mathcal{B}(r)\mathcal{C}(r)]$. This formulation mirrors the classical motion of a particle with unit mass under the influence of a potential $V(r)$. The photon’s motion is therefore constrained to regions where $V(r)\geq 0$. The asymptotic behavior of the spacetime ensures that, as $r\to\infty$, the potential tends toward a positive constant: $V(r)\to E^{2}>0$, indicating that photons can exist at arbitrarily large distances.
We concentrate on the lensing configuration where a photon, emitted from a distant source, travels toward the compact object, reaches a minimal radial coordinate $r_{0}$, and then escapes back to spatial infinity. For such a deflection process to take place rather than being trapped in a circular orbit, it is required that $r_{0}>r_{ph}$. The turning point of the trajectory, $r=r_{0}$, is characterized as the outermost positive root of $R(r)=0$, at which the effective potential satisfies $V(r_{0})=0$. Provided that the functions $\mathcal{B}(r)$ and $\mathcal{C}(r)$ are regular at this point, one can infer from the definition of $R(r)$ the following condition at the point of closest approach:
$\mathcal{A}_{0}\dot{t}^{2}_{0}=\mathcal{C}_{0}\dot{\phi}^{2}_{0}.$ (12)
It is worth mentioning that any quantity carrying the subscript ”00” is understood to be evaluated at the radial distance $r=r_{0}$, which represents the point of closest approach of the photon to the lensing object. Since we are interested in the trajectory of a single light ray, we take the impact parameter $b$ to be positive—an assumption that does not affect the generality of the treatment. Given that $b$ is a positive quantity along the path of the photon, it can be determined directly at $r_{0}$ by the following relation:
$b(r_{0})=\frac{L}{E}=\frac{\mathcal{C}_{0}\dot{\phi}_{0}}{\mathcal{A}_{0}\dot{t}_{0}}=\sqrt{\frac{\mathcal{C}_{0}}{\mathcal{A}_{0}}}.$ (13)
It is also worth noting that the function $R(r)$ can alternatively be written in the equivalent form:
$R(r)=\frac{\mathcal{A}_{0}\mathcal{C}}{\mathcal{A}\mathcal{C}_{0}}-1.$ (14)
Following the methodology introduced in Ref. hasse2002gravitational , one can establish a criterion that fully characterizes the presence of circular photon orbits. This condition is not only required but also guarantees the existence of such trajectories. Based on this framework, the corresponding equation governing the path of light takes the form:
$\frac{\mathcal{B}\mathcal{C}\dot{r}^{2}}{E^{2}}+b^{2}=\frac{\mathcal{C}}{\mathcal{A}},$ (15)
so that
$\ddot{r}+\frac{1}{2}\left(\frac{\mathcal{B}^{\prime}}{\mathcal{B}}+\frac{\mathcal{C}^{\prime}}{\mathcal{C}}\dot{r}^{2}\right)=\frac{E^{2}\mathcal{D}}{\mathcal{A}\mathcal{B}}.$ (16)
In the region where $r\geq r_{ph}$, the metric functions $\mathcal{A}(r)$, $\mathcal{B}(r)$, and $\mathcal{C}(r)$ are assumed to be regular, strictly positive, and finite. Given a positive energy $E$, the condition $\mathcal{D}(r)=0$ serves as the fundamental requirement for the existence of a circular photon orbit. At the radius of the photon sphere, $r=r_{ph}$, we find that the derivative of $R(r)$ vanishes, expressed as $R^{\prime}_{ph}=\mathcal{D}_{ph}\mathcal{C}_{ph}\mathcal{A}_{ph}/b^{2}=0$. Here, the subscript “$ph$” indicates evaluation at the photon sphere.
We now introduce the critical value of the impact parameter, denoted as $b_{c}$, which is
$b_{c}(r_{ph})\equiv\lim_{r_{0}\to r_{ph}}\sqrt{\frac{\mathcal{C}_{0}}{\mathcal{A}_{0}}}.$ (17)
From this point onward, the scenario under consideration will be classified as the strong deflection regime. To proceed further, we compute the radial derivative of the effective potential $V(r)$, which yields:
$V^{\prime}(r)=\frac{L^{2}}{\mathcal{B}\mathcal{C}}\left[R^{\prime}+\left(\frac{\mathcal{C}^{\prime}}{\mathcal{C}}-\frac{\mathcal{B}^{\prime}}{\mathcal{B}}\right)R\right].$ (18)
In the regime of strong gravitational deflection, as the point of closest approach $r_{0}$ approaches the photon sphere radius $r_{ph}$, both the effective potential $V(r_{0})$ and its radial derivative $V^{\prime}(r_{0})$ simultaneously vanish. This limiting behavior reshapes the equation governing the light’s path into the simplified expression:
$\left(\frac{\mathrm{d}r}{\mathrm{d}\phi}\right)^{2}=\frac{R(r)\mathcal{C}(r)}{\mathcal{B}(r)}.$ (19)
Accordingly, the quantity known as the deflection angle, denoted by $\alpha(r_{0})$, takes the following integral form in terms of the closest approach distance $r_{0}$:
$\alpha(r_{0})=I(r_{0})-\pi,$ (20)
with the function $I(r_{0})$, which encodes the integral contribution to the deflection angle, is defined as follows:
$I(r_{0})\equiv 2\int^{\infty}_{r_{0}}\frac{\mathrm{d}r}{\sqrt{\frac{R(r)\mathcal{C}(r)}{\mathcal{B}(r)}}}.$ (21)
As a first step in the analysis, attention must be turned to the integral expression governing the deflection angle. However, as emphasized in Tsukamoto’s investigation tsukamoto2017deflection , evaluating this integral analytically is far from straightforward due to the divergence that arises near the photon sphere. To facilitate the computation, we introduce the following auxiliary function, as also suggested in Ref. tsukamoto2017deflection :
$z\equiv 1-\frac{r_{0}}{r}.$ (22)
Here, with this new definition in place, the integral expression can be reformulated in the following manner:
$I(r_{0})=\int^{1}_{0}f(z,r_{0})\mathrm{d}z,$ (23)
where
$f(z,z_{0})\equiv\frac{2r_{0}}{\sqrt{G(z,r_{0})}},\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,G(z,r_{0})\equiv R\frac{\mathcal{C}}{\mathcal{B}}(1-z)^{4}.$ (24)
Expressed in terms of the variable $z$, the function $R(r)$ can be rewritten as:
$R(r)=\mathcal{D}_{0}r_{0}z+\left[\frac{r_{0}}{2}\left(\frac{\mathcal{C}^{\prime\prime}_{0}}{\mathcal{C}_{0}}-\frac{\mathcal{A}_{0}^{\prime\prime}}{\mathcal{A}_{0}}\right)+\left(1-\frac{\mathcal{A}_{0}^{\prime}r_{0}}{\mathcal{A}_{0}}\right)D_{0}\right]r_{0}z^{2}+\mathcal{O}(z^{3})+...\,\,\,\,.$ (25)
We now proceed by expanding the function $G(z,r_{0})$ in a power series around $z=0$. This yields:
$G(z,r_{0})=\sum^{\infty}_{n=1}c_{n}(r_{0})z^{n},$ (26)
in which the quantities $c_{1}(r)$ and $c_{2}(r)$ are expressed as
$c_{1}(r_{0})=\frac{\mathcal{C}_{0}\mathcal{D}_{0}r_{0}}{\mathcal{B}_{0}},$ (27)
and
$c_{2}(r_{0})=\frac{\mathcal{C}_{0}r_{0}}{\mathcal{B}_{0}}\left\{\mathcal{D}_{0}\left[\left(\mathcal{D}_{0}-\frac{\mathcal{B}^{\prime}_{0}}{\mathcal{B}_{0}}\right)r_{0}-3\right]+\frac{r_{0}}{2}\left(\frac{\mathcal{C}^{\prime\prime}_{0}}{\mathcal{C}_{0}}-\frac{\mathcal{A}^{\prime\prime}_{0}}{\mathcal{A}_{0}}\right)\right\}.$ (28)
Additionally, in the regime of strong gravitational deflection, one finds that
$c_{1}(r_{ph})=0,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,c_{2}(r_{ph})=\frac{\mathcal{C}_{ph}r^{2}_{ph}}{2\mathcal{B}_{ph}}\mathcal{D}^{\prime}_{ph},\,\,\,\,\,\,\,\text{with}\,\,\,\,\,\mathcal{D}^{\prime}_{ph}=\frac{\mathcal{C}^{\prime\prime}}{\mathcal{C}_{ph}}-\frac{\mathcal{A}^{\prime\prime}}{\mathcal{A}_{ph}}.$ (29)
In this context, $G(z,r_{0})$ is written in a more compact form, as follows:
$G_{ph}(z)=c_{2}(r_{ph})z^{2}+\mathcal{O}(z^{3}).$ (30)
The divergence in the function $f(z,r_{0})$ becomes evident through its leading behavior, which scales as $1/z$ near $z=0$. This singularity gives rise to a logarithmic divergence in the integral $I(r_{0})$ as the closest approach radius $r_{0}$ approaches the photon sphere radius $r_{ph}$. To handle this issue analytically, the integral $I(r_{0})$ is decomposed into two separate contributions: a singular term, denoted by $I_{D}(r_{0})$, which takes into account the divergent structure, and a finite remainder, referred to as $I_{R}(r_{0})$. Accordingly, the divergent component takes the form:
$I_{D}(r_{0})\equiv\int^{1}_{0}f_{D}(z,r_{0})\mathrm{d}z,\,\,\,\,\,\,\,\text{with}\,\,\,\,\,\,f_{D}(z,r_{0})\equiv\frac{2r_{0}}{\sqrt{c_{1}(r_{0})z+c_{2}(r_{0})z^{2}}}.$ (31)
Upon performing the integration, the result becomes
$I_{D}(r_{0})=\frac{4r_{0}}{\sqrt{c_{2}(r_{0})}}\ln\left[\frac{\sqrt{c_{2}(r_{0})}+\sqrt{c_{1}(r_{0})+c_{2}(r_{0})}}{\sqrt{c_{1}(r_{0})}}\right].$ (32)
Expanding $c_{1}(r_{0})$ and $b(r_{0})$ in a series around the point $r_{0}=r_{ph}$
$c_{1}(r_{0})=\frac{\mathcal{C}_{ph}r_{ph}\mathcal{D}^{\prime}_{ph}}{\mathcal{B}_{ph}}(r_{0}-r_{ph})+\mathcal{O}((r_{0}-r_{ph})^{2}),$ (33)
and
$b(r_{0})=b_{c}(r_{ph})+\frac{1}{4}\sqrt{\frac{\mathcal{C}_{ph}}{\mathcal{A}_{ph}}}\mathcal{D}^{\prime}_{ph}(r_{0}-r_{ph})^{2}+\mathcal{O}((r_{0}-r_{ph})^{3}),$ (34)
which yields the following expression in the strong deflection regime
$\lim_{r_{0}\to r_{ph}}c_{1}(r_{0})=\lim_{b\to b_{c}}\frac{2\mathcal{C}_{ph}r_{ph}\sqrt{\mathcal{D}^{\prime}}}{B_{ph}}\left(\frac{b}{b_{c}}-1\right)^{1/2}.$ (35)
In this way, we write $I_{D}(b)$ as being
$I_{D}(b)=-\frac{r_{ph}}{\sqrt{c_{2}(r_{ph})}}\ln\left[\frac{b}{b_{c}}-1\right]+\frac{r_{ph}}{\sqrt{c_{2}(r_{ph})}}\ln\left[r^{2}\mathcal{D}^{\prime}_{ph}\right]+\mathcal{O}[(b-b_{c})\ln(b-b_{c})].$ (36)
Moreover, the regular contribution $I_{R}(b)$ is defined as
$I_{R}(b)=\int^{0}_{1}f_{R}(z,b_{c})\mathrm{d}z+\mathcal{O}[(b-b_{c})\ln(b-b_{c})].$ (37)
Let us define the regular part of the integrand as $f_{R}=f(z,r_{0})-f_{D}(z,r_{0})$. Under the strong deflection approximation, the total deflection angle then takes the form
$a(b)=-\tilde{a}\ln\left[\frac{b}{b_{c}}-1\right]+\tilde{b}+\mathcal{O}[(b-b_{c})\ln(b-b_{c})],$ (38)
with
$\tilde{a}=\sqrt{\frac{2\mathcal{B}_{ph}A_{ph}}{\mathcal{C}^{\prime\prime}_{ph}\mathcal{A}_{ph}-\mathcal{C}_{ph}\mathcal{A}^{\prime\prime}_{ph}}},\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,\tilde{b}=\tilde{a}\ln\left[r^{2}_{ph}\left(\frac{\mathcal{C}^{\prime\prime}}{\mathcal{C}_{ph}}-\frac{\mathcal{A}^{\prime\prime}_{ph}}{\mathcal{C}_{ph}}\right)\right]+I_{R}(r_{ph})-\pi.$ (39)